# On the Robustness of Kernel Goodness-of-Fit Tests

### v4

@@ A.3.3 Proof of Lemma˜13
Under the assumed conditions on $k$ and assuming $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$, Gorham and Mackey, (2017, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$, so the V-statistic $D^{2}(\mathbb{X}_{n}^{\ast})$ is degenerate (see, e.g., Serfling,, 2009, Chapter 6), and classic results on the asymptotics of degenerate V-statistics (Serfling,, 2009, Theorem 6.4.1 A) shows that $nD^{2}(\mathbb{X}_{n}^{\ast})$ converges weakly to a non-negative distribution. Since the square-root function $x\mapsto\sqrt{x}$ is continuous on $0,\infty)$ and continuous functions preserve weak limits by the Continuous Mapping Theorem (Van der Vaart,, 2000, Theorem 2.3), the squared-root statistic, $\sqrt{n}D(\mathbb{X}_{n}^{\ast})$, also converges weakly to a non-negative distribution. In particular, the scaled quantile $\sqrt{n}q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})$, and thus also $\sqrt{n}\zeta_{n,\theta}$, converge to a positive number as $n\to\infty$. Also since $\eta\left(f_{n}\right)=o(n^{-1/2})$ as argued below (43), we have $\sqrt{n}\eta\left(f_{n}\right)\to 0$. In summary, we have shown that
@@ A.3.4 Proofs for Proposition˜14 and Theorem˜3
To show $\mathcal{B}$ occurs with high probability, we use a concentration inequality for Binomial distributions; see, e.g., Chung and Lu, (2002, Lemma 2.1, Eq. (2.2)). Applying this result to $\text{Binomial}(n,\epsilon_{n})$, we have
@@ A.4 $P$-Targeted Kernel Stein Discrepancy
This section states preliminary results which will be needed to prove Theorem˜4. The main tool we use is the following generalized definition of KSD, which was originally proposed in Shi and Mackey, (2024, Definition 4).
@@
It can be shown that $P$-KSD is a Maximum Mean Discrepancy (MMD, Fortet and Mourier,, 1953; Müller,, 1997) with the Stein reproducing kernel $u_{p}$ (Barp et al.,, 2024, Theorem 1). When either of the two arguments coincides with $P$, the $P$-KSD reduces to the standard KSD, as we will show in the next section. The main benefit of working with $P$-KSD rather than KSD is that $P$-KSD satisfies both symmetry and a triangle inequality (Shi and Mackey,, 2024, Lemma 1), i.e., $\mathbb{S}_{P}(Q,R)=\mathbb{S}_{P}(R,Q)$ and $\mathbb{S}_{P}(Q,R)\leq\mathbb{S}_{P}(Q,R^{\prime})+\mathbb{S}_{P}(R^{\prime},R)$, for any $Q,R,R^{\prime}\in\mathcal{P}(\mathbb{R}^{d})$ for which they are defined. These properties allow us to show several lemmas, which will be crucial in proving the validity of our robust-KSD tests in later sections.
@@
  • Lemma˜18 restates Tolstikhin et al., (2017, Proposition A.1), which is a McDiarmid-type inequality for MMD. It also applies to $P$-KSD, as an $P$-KSD is also an MMD. We will use it in Appendix˜D to derive a robust-KSD test that is well-calibrated for finite samples. Other deviation bounds instead of the McDiarmid bound could also be used to construct a similar test;see Remark˜24.1 for a discussion.
The MMD has already been studied extensively in the context of both Bayesian and frequentist robust parameter estimation; see, e.g., Briol et al., (2019); Chérief-Abdellatif and Alquier, (2020, 2022); Dellaporta et al., (2022); Dellaporta and Damoulas, (2023); Alquier and Gerber, (2024). A key assumption to ensure robustness in all of these papers is that the kernel is assumed to be bounded, and this is exactly what we are able to achieve with the KSD and $P$-KSD through Lemma˜2.
@@ A.4.1 Properties of $P$-KSD
Proof of Lemma˜15  This follows directly from the symmetry of $P$-KSD (Shi and Mackey,, 2024, Lemma 1) and that the assumed conditions imply $\mathbb{E}_{\mathbf{X}\sim P}[(\mathcal{A}_{P}f)(\mathbf{X})]=0$ for all $f\in\mathcal{B}$ (Gorham and Mackey,, 2017, Proposition 1), so one of the two expectations in (48) vanishes when either of the two arguments coincides with $P$.
Barp et al., (2024, Theorem 1) shows that a KSD is equivalent to an MMD with the Stein kernel $u_{p}$. Since $P$-KSD is a KSD by Lemma˜15, we can conclude that the $P$-KSD is also an MMD with kernel $u_{p}$, and that $P$-KSD has a double-expectation form similar to MMD.
@@
Proof of Lemma˜16  The assumed conditions ensure $\mathbb{S}_{P}^{2}(Q,R)$ is well-defined. Since $P$-KSD is an MMD with reproducing kernel $u_{p}$, we can apply Gretton et al., (2012, Lemma 4) to conclude the first equality, and Gretton et al., (2012, Lemma 6) to yield the second.
@@
Proof of Lemma˜17  The $P$-KSD $\mathbb{S}_{P}(R,P)$ is well-defined under the assumed conditions, and hence so is $\mathbb{S}_{P}\big{(}\lambda R+(1-\lambda)P,P\big{)}$ for any $\lambda\in[0,1]$. Also since $P$-KSD is an MMD with (potentially unbounded) reproducing kernel $u_{p}$, the claim then follows by Sun and Zou, (2023, Proposition 3) and noting that their proof extends directly to potentially unbounded kernels.
@@
Proof of Lemma˜18  Let $Q_{n}$ denote the empirical distribution based on $\mathbb{X}_{n}$. By the first equality in Lemma˜16, we can write $\mathbb{S}_{P}(\mathbb{X}_{n},Q)=\|\mathbb{E}_{\mathbf{X}\sim Q_{n}}[u_{p}(\cdot,\mathbf{X})]-\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{X})]\|_{\mathcal{H}_{u}}$. Under the assumptions in Lemma˜2, the function $u_{p}(\cdot,\mathbf{x}):\mathbb{R}^{d}\to\mathcal{H}_{u}$ is continuous for all $\mathbf{x}\in\mathbb{R}^{d}$. Moreover, $\|u_{p}(\cdot,\mathbf{x})\|_{\mathcal{H}_{u}}=u_{p}(\mathbf{x},\mathbf{x})$ by the reproducing property of the Stein kernel $u_{p}$ (see also the argument before (30)), and $\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})<\infty$ again by Lemma˜2. We can hence apply the McDiarmid-type inequality for MMD in Tolstikhin et al., (2017, Proposition A.1) to conclude the claimed result.
@@ A.5 Validity of the Bootstrap Approach
The bootstrap procedure used in robust-KSD is the same as the one in the standard KSD test, and it is not immediately obvious that this gives a valid decision threshold. For the standard KSD test, the V-statistic $D^{2}(\mathbb{X}_{n})$ is degenerate under the point null $H_{0}:Q=P$, i.e., $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,P)]=0$ when $Q=P$ (Liu et al.,, 2016, Theorem 4.1), and classic bootstrapping methods for degenerate V-statistics (Arcones and Gine,, 1992; Huskova and Janssen,, 1993) show that the bootstrap quantile $q_{\infty,1-\alpha}$ is a valid decision threshold. However, the same argument cannot be applied directly to our robust test, because our null $H^{\mathrm{C}}_{0}$ is a composite one that contains not only $P$ but also other distributions $Q\neq P$, in which case $D^{2}(\mathbb{X}_{n})$ is no longer degenerate as shown by Liu et al., (2016, Theorem 4.1). Nevertheless, our proof shows that $q_{\infty,1-\alpha}$ is still a correct decision threshold.
Intuition of the proof. Let $\mathbb{X}_{n}=\{\mathbf{X}_{i}\}_{i=1}^{n}$ be a realisations from some $Q\in\mathcal{P}(\mathbb{R}^{d})$. We will first show that the rejection probability of the robust-KSD test can be bounded as
@@
Our key observation is that $\mathbb{S}_{P}^{2}(\mathbb{X}_{n},Q)$ is the second-order term in the Hoeffding’s decomposition (Arcones and Gine,, 1992, Eq. 3.2) of the V-statistic estimate $D^{2}(\mathbb{X}_{n},P)$ for the standard KSD defined in (2). This suggests that (3) is an appropriate bootstrap sample. Crucially, this argument does not require $Q=P$, unlike in the proof of the standard KSD test. We formalize this argument in the next section.
A naive alternative approach to bootstrap $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ is to note that it is degenerate for any $Q$ and use standard bootstrapping techniques for degenerate V-statistics (e.g., Arcones and Gine,, 1992, Theorem 3.5). However, this approach is not applicable here since, to compute the bootstrap sample, it requires the function $u_{p}(\mathbf{x},\mathbf{x}^{\prime};Q)$ to be computable at any $\mathbf{x},\mathbf{x}^{\prime}$, which is not the case since $u_{p}(\mathbf{x},\mathbf{x}^{\prime};Q)$ involves intractable expectations over $Q$ (recall that $u_{p}$ has mean zero under $P$, but not necessarily under $Q$).
@@ A.5.1 Bootstrapping the $P$-KSD Estimate
Our proof for Theorem˜4 relies on Arcones and Gine, (1992, Lemma 3.4), which states that the asymptotic distribution of the second-order term in the Hoeffding’s decomposition of a V-statistic can be approximated by a bootstrap distribution. We restate this result for the case of $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ in Lemma˜19.
Lemma 19 (Arcones and Gine, (1992), Lemma 3.4).
@@
Proof of Lemma˜20  As argued in (A.5), the squared $P$-KSD $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ can be written as a V-statistic with symmetric function $u_{p}(\cdot,\cdot;Q)$. Moreover, direct computation shows that $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{X};Q)]\equiv 0$, so the symmetric function $u_{p}(\cdot,\cdot;Q)$ is $Q$-degenerate of order 1 (Arcones and Gine,, 1992, pp. 5). On the other hand, it is well-known that the weighted bootstrap form can be equivalently written as an Efron’s resampled bootstrap statistic (Janssen,, 1994; Dehling and Mikosch,, 1994; Janssen,, 1997), i.e.,
@@
The claim then follows by noting that the mapping $u\mapsto\sqrt{u}$ is everywhere continuous on $0,\infty)$ and that weak convergence is preserved by continuous function by the Continuous Mapping Theorem (Van der Vaart,, 2000, Theorem 2.3).
@@ A.5.2 Proof of Theorem˜4
$\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]$
$\leq\;\mathbb{E}_{\mathbf{X}\sim Q}\Big{[}\big{(}\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}^{2}h(0)+2\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}\|\nabla w(\mathbf{x})\|_{2}h(0)+\|\nabla w(\mathbf{x})\|_{2}^{2}h(0)+w(\mathbf{x})^{2}|\nabla^{\top}\nabla h(0)|\big{)}^{2}\Big{]}$
 $\leq\;4\Big{(}\mathbb{E}_{\mathbf{X}\sim Q}\big{[}\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}^{4}\big{]}h^{2}(0)+2\mathbb{E}_{\mathbf{X}\sim Q}\big{[}\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}^{2}\|\nabla w(\mathbf{x})\|_{2}^{2}\big{]}h^{2}(0)+\mathbb{E}_{\mathbf{X}\sim Q}\big{[}\|\nabla w(\mathbf{x})\|_{2}^{4}\big{]}h^{2}(0)$
$\quad+\mathbb{E}_{\mathbf{X}\sim Q}\big{[}w(\mathbf{x})^{4}\big{]}|\nabla^{\top}\nabla h(0)|^{2}\Big{)}\;,$
@@
where (53) holds since it can be checked that $\Delta_{\theta}(\mathbb{X}_{n})>\gamma$ if and only if $D(\mathbb{X}_{n})-\theta>\gamma$ for any $\gamma\geq 0$. The argument in Liu et al., (2016, Theorem 4.1) shows that that $\sqrt{n}(D^{2}(\mathbb{X}_{n})-D^{2}(Q,P))$ converges weakly to a Gaussian limit assuming $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})^{2}]<\infty$, which holds since (30) and Jensen’s inequality imply
@@
Remark 20.1.
If the bootstrapped quantile $q_{\infty,1-\alpha}(\mathbb{X}_{n})$ is replaced by the quantile $q^{\ast}_{1-\alpha}(\mathbb{X}_{n})$ of the distribution of the $P$-KSD $\mathbb{S}_{P}(\mathbb{X}_{n},Q)$, then the bootstrap approximation error in (52) vanishes, i.e., $\delta_{n}=0$ for all $n$. We can then prove the following stronger, uniform level control

### v5

@@ A.3.3 Proof of Lemma˜13
Under the assumed conditions on $k$ and assuming $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$, Gorham and Mackey (2017, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$, so the V-statistic $D^{2}(\mathbb{X}_{n}^{\ast})$ is degenerate (see, e.g., Serfling, 2009, Chapter 6), and classic results on the asymptotics of degenerate V-statistics (Serfling, 2009, Theorem 6.4.1 A) shows that $nD^{2}(\mathbb{X}_{n}^{\ast})$ converges weakly to a non-negative distribution. Since the square-root function $x\mapsto\sqrt{x}$ is continuous on $0,\infty)$ and continuous functions preserve weak limits by the Continuous Mapping Theorem (Van der Vaart, 2000, Theorem 2.3), the squared-root statistic, $\sqrt{n}D(\mathbb{X}_{n}^{\ast})$, also converges weakly to a non-negative distribution. In particular, the scaled quantile $\sqrt{n}q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})$, and thus also $\sqrt{n}\zeta_{n,\theta}$, converge to a positive number as $n\to\infty$. Also since $\eta\left(f_{n}\right)=o(n^{-1/2})$ as argued below (43), we have $\sqrt{n}\eta\left(f_{n}\right)\to 0$. In summary, we have shown that
@@ A.3.4 Proofs for Proposition˜14 and Theorem˜3
To show $\mathcal{B}$ occurs with high probability, we use a concentration inequality for Binomial distributions; see, e.g., Chung and Lu (2002, Lemma 2.1, Eq. 2.2). Applying this result to $\text{Binomial}(n,\epsilon_{n})$, we have
@@ A.4 $P$-Targeted Kernel Stein Discrepancy
This section states preliminary results which will be needed to prove Theorem˜4. The main tool we use is the following generalized definition of KSD, which was originally proposed in Shi and Mackey (2024, Definition 4).
@@
It can be shown that $P$-KSD is a Maximum Mean Discrepancy (MMD, Fortet and Mourier, 1953; Müller, 1997) with the Stein reproducing kernel $u_{p}$ (Barp et al., 2024, Theorem 1). When either of the two arguments coincides with $P$, the $P$-KSD reduces to the standard KSD, as we will show in the next section. The main benefit of working with $P$-KSD rather than KSD is that $P$-KSD satisfies both symmetry and a triangle inequality (Shi and Mackey, 2024, Lemma 1), i.e., $\mathbb{S}_{P}(Q,R)=\mathbb{S}_{P}(R,Q)$ and $\mathbb{S}_{P}(Q,R)\leq\mathbb{S}_{P}(Q,R^{\prime})+\mathbb{S}_{P}(R^{\prime},R)$, for any $Q,R,R^{\prime}\in\mathcal{P}(\mathbb{R}^{d})$ for which they are defined. These properties allow us to show several lemmas, which will be crucial in proving the validity of our robust-KSD tests in later sections.
@@
  • Lemma˜18 restates Tolstikhin et al. (2017, Proposition A.1), which is a McDiarmid-type inequality for MMD. It also applies to $P$-KSD, as an $P$-KSD is also an MMD. We will use it in Appendix˜D to derive a robust-KSD test that is well-calibrated for finite samples. Other deviation bounds instead of the McDiarmid bound could also be used to construct a similar test;see Remark˜10 for a discussion.
The MMD has already been studied extensively in the context of both Bayesian and frequentist robust parameter estimation; see, e.g., Briol et al. (2019); Chérief-Abdellatif and Alquier (2020); Chérief-Abdellatif and Alquier (2022); Dellaporta et al. (2022); Dellaporta and Damoulas (2023); Alquier and Gerber (2024). A key assumption to ensure robustness in all of these papers is that the kernel is assumed to be bounded, and this is exactly what we are able to achieve with the KSD and $P$-KSD through Lemma˜2.
@@ A.4.1 Properties of $P$-KSD
Proof of Lemma˜15 This follows directly from the symmetry of $P$-KSD (Shi and Mackey, 2024, Lemma 1) and that the assumed conditions imply $\mathbb{E}_{\mathbf{X}\sim P}[(\mathcal{A}_{P}f)(\mathbf{X})]=0$ for all $f\in\mathcal{B}$ (Gorham and Mackey, 2017, Proposition 1), so one of the two expectations in (48) vanishes when either of the two arguments coincides with $P$.
Barp et al. (2024, Theorem 1) shows that a KSD is equivalent to an MMD with the Stein kernel $u_{p}$. Since $P$-KSD is a KSD by Lemma˜15, we can conclude that the $P$-KSD is also an MMD with kernel $u_{p}$, and that $P$-KSD has a double-expectation form similar to MMD.
@@
Proof of Lemma˜16 The assumed conditions ensure $\mathbb{S}_{P}^{2}(Q,R)$ is well-defined. Since $P$-KSD is an MMD with reproducing kernel $u_{p}$, we can apply Gretton et al. (2012, Lemma 4) to conclude the first equality, and Gretton et al. (2012, Lemma 6) to yield the second.
@@
Proof of Lemma˜17 The $P$-KSD $\mathbb{S}_{P}(R,P)$ is well-defined under the assumed conditions, and hence so is $\mathbb{S}_{P}\big(\lambda R+(1-\lambda)P,P\big)$ for any $\lambda\in[0,1]$. Also since $P$-KSD is an MMD with (potentially unbounded) reproducing kernel $u_{p}$, the claim then follows by Sun and Zou (2023, Proposition 3) and noting that their proof extends directly to potentially unbounded kernels.
@@
Proof of Lemma˜18 Let $Q_{n}$ denote the empirical distribution based on $\mathbb{X}_{n}$. By the first equality in Lemma˜16, we can write $\mathbb{S}_{P}(\mathbb{X}_{n},Q)=\|\mathbb{E}_{\mathbf{X}\sim Q_{n}}[u_{p}(\cdot,\mathbf{X})]-\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{X})]\|_{\mathcal{H}_{u}}$. Under the assumptions in Lemma˜2, the function $u_{p}(\cdot,\mathbf{x}):\mathbb{R}^{d}\to\mathcal{H}_{u}$ is continuous for all $\mathbf{x}\in\mathbb{R}^{d}$. Moreover, $\|u_{p}(\cdot,\mathbf{x})\|_{\mathcal{H}_{u}}=u_{p}(\mathbf{x},\mathbf{x})$ by the reproducing property of the Stein kernel $u_{p}$ (see also the argument before (30)), and $\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})<\infty$ again by Lemma˜2. We can hence apply the McDiarmid-type inequality for MMD in Tolstikhin et al. (2017, Proposition A.1) to conclude the claimed result.
@@ A.5 Validity of the Bootstrap Approach
The bootstrap procedure used in robust-KSD is the same as the one in the standard KSD test, and it is not immediately obvious that this gives a valid decision threshold. For the standard KSD test, the V-statistic $D^{2}(\mathbb{X}_{n})$ is degenerate under the point null $H_{0}:Q=P$, i.e., $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,P)]=0$ when $Q=P$ (Liu et al., 2016, Theorem 4.1), and classic bootstrapping methods for degenerate V-statistics (Arcones and Gine, 1992; Huskova and Janssen, 1993) show that the bootstrap quantile $q_{\infty,1-\alpha}$ is a valid decision threshold. However, the same argument cannot be applied directly to our robust test, because our null $H^{\mathrm{C}}_{0}$ is a composite one that contains not only $P$ but also other distributions $Q\neq P$, in which case $D^{2}(\mathbb{X}_{n})$ is no longer degenerate as shown by Liu et al. (2016, Theorem 4.1). Nevertheless, our proof shows that $q_{\infty,1-\alpha}$ is still a correct decision threshold.
Intuition of the proof. Let $\mathbb{X}_{n}=\{\mathbf{X}_{i}\}_{i=1}^{n}$ be a random sample from some $Q\in\mathcal{P}(\mathbb{R}^{d})$. We will first show that the rejection probability of the robust-KSD test can be bounded as
@@
Our key observation is that $\mathbb{S}_{P}^{2}(\mathbb{X}_{n},Q)$ is the second-order term in the Hoeffding’s decomposition (Arcones and Gine, 1992, Eq. 3.2) of the V-statistic estimate $D^{2}(\mathbb{X}_{n},P)$ for the standard KSD defined in (2). This suggests that (3) is an appropriate bootstrap sample. Crucially, this argument does not require $Q=P$, unlike in the proof of the standard KSD test. We formalize this argument in the next section.
A naive alternative approach to bootstrap $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ is to note that it is degenerate for any $Q$ and use standard bootstrapping techniques for degenerate V-statistics (e.g., Arcones and Gine, 1992, Theorem 3.5). However, this approach is not applicable here since, to compute the bootstrap sample, it requires the function $u_{p}(\mathbf{x},\mathbf{x}^{\prime};Q)$ to be computable at any $\mathbf{x},\mathbf{x}^{\prime}$, which is not the case since $u_{p}(\mathbf{x},\mathbf{x}^{\prime};Q)$ involves intractable expectations over $Q$ (recall that $u_{p}$ has mean zero under $P$, but not necessarily under $Q$).
@@ A.5.1 Bootstrapping the $P$-KSD Estimate
Our proof for Theorem˜4 relies on Arcones and Gine (1992, Lemma 3.4), which states that the asymptotic distribution of the second-order term in the Hoeffding’s decomposition of a V-statistic can be approximated by a bootstrap distribution. We restate this result for the case of $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ in Lemma˜19.
Lemma 19 (Arcones and Gine (1992), Lemma 3.4).
@@
Proof of Lemma˜20 As argued in (49), the squared $P$-KSD $\mathbb{S}^{2}_{P}(\mathbb{X}_{n},Q)$ can be written as a V-statistic with symmetric function $u_{p}(\cdot,\cdot;Q)$. Moreover, direct computation shows that $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{X};Q)]\equiv 0$, so the symmetric function $u_{p}(\cdot,\cdot;Q)$ is $Q$-degenerate of order 1 (Arcones and Gine, 1992, pp. 5). On the other hand, it is well-known that the weighted bootstrap form can be equivalently written as an Efron’s resampled bootstrap statistic (Janssen, 1994; Dehling and Mikosch, 1994; Janssen, 1997), i.e.,
@@
The claim then follows by noting that the mapping $u\mapsto\sqrt{u}$ is everywhere continuous on $0,\infty)$ and that weak convergence is preserved by continuous function by the Continuous Mapping Theorem (Van der Vaart, 2000, Theorem 2.3).
@@ A.5.2 Proof of Theorem˜4
$\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]\;$ $\leq\;\mathbb{E}_{\mathbf{X}\sim Q}\bigg[\Big(\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}^{2}h(0)+2\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}\|\nabla w(\mathbf{x})\|_{2}h(0)+\|\nabla w(\mathbf{x})\|_{2}^{2}h(0)$
$\;\qquad\qquad\quad+w(\mathbf{x})^{2}\big|\nabla^{\top}\nabla h(0)\big|\Big)^{2}\bigg]$
 $\leq\;4\Big(\mathbb{E}_{\mathbf{X}\sim Q}\big[\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}^{4}\big]h^{2}(0)+2\mathbb{E}_{\mathbf{X}\sim Q}\big[\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}^{2}\|\nabla w(\mathbf{x})\|_{2}^{2}\big]h^{2}(0)$
$\quad\quad\;\;+\mathbb{E}_{\mathbf{X}\sim Q}\big[\|\nabla w(\mathbf{x})\|_{2}^{4}\big]h^{2}(0)+\mathbb{E}_{\mathbf{X}\sim Q}\big[w(\mathbf{x})^{4}\big]\big|\nabla^{\top}\nabla h(0)\big|^{2}\Big)\;,$
@@
where (53) holds since it can be checked that $\Delta_{\theta}(\mathbb{X}_{n})>\gamma$ if and only if $D(\mathbb{X}_{n})-\theta>\gamma$ for any $\gamma\geq 0$. The argument in Liu et al. (2016, Theorem 4.1) shows that that $\sqrt{n}(D^{2}(\mathbb{X}_{n})-D^{2}(Q,P))$ converges weakly to a Gaussian limit assuming $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})^{2}]<\infty$, which holds since (30) and Jensen’s inequality imply
@@
Remark 9.
If the bootstrapped quantile $q_{\infty,1-\alpha}(\mathbb{X}_{n})$ is replaced by the quantile $q^{\ast}_{1-\alpha}(\mathbb{X}_{n})$ of the distribution of the $P$-KSD $\mathbb{S}_{P}(\mathbb{X}_{n},Q)$, then the bootstrap approximation error in (52) vanishes, i.e., $\delta_{n}=0$ for all $n$. In that case, we have the following stronger, uniform level control